OpenStudy (anonymous):

Need help: a2 + 12ab + 27b2 Factoring!

OpenStudy (anonymous):

\[a ^{2} + 12ab + 27b ^{2}\]

hartnn (hartnn):

can you find 2 numbers with product = 27 and sum =12 ?

OpenStudy (anonymous):

No?

OpenStudy (anonymous):

Sorry, I asked the wrong question. I'll post the correct one.

OpenStudy (anonymous):

*facepalm* yes 3 & 9

hartnn (hartnn):

good.

hartnn (hartnn):

so, split 12ab = 3ab and 9ab

hartnn (hartnn):

\(a^2+3ab+9ab+27b^2 \) factor out a from 1st 2 terms and 9b from last two, what u get ?

OpenStudy (anonymous):

This is what I have so far . . . a2 + 12ab + 27b^2 3*9 = 27 3 + 9 = 12 (ab + 3) (ab + 9)??

hartnn (hartnn):

\(a^2+3ab+9ab+27b^2 = a(a+3b)+9b(a+3b)\\= (a+3b)(a+9b)\) got this ?

OpenStudy (anonymous):

ahhh ok